OpenStudy (anonymous):

Wayne is buying a house for $205,000. He is financing $175,000 and obtained a 30-year, fixed-rate mortgage with a 6.425% interest rate. How much are his monthly payments? $1,285.64 $936.48 $1,517.23 $1,097.50

OpenStudy (anonymous):

what formula does your book give for the monthly payments?

OpenStudy (anonymous):

M=P x i(1+i)^nt / (1+i)^nt -1

OpenStudy (anonymous):

so P = 175000 (amount financed) t = 30 (period of loan in years) n = 12 ( number of payments per year) i = 6.425/100 = 0.06425 (APR) \[ M=175000\times\frac{0.06425\times(1+0.06425)^{12\times30}}{(1+0.06425)^{12\times30}-1} \]

OpenStudy (anonymous):

here, I missed to divide the APR by "12"!!!